The displacement (in meters) of a particle executing simple harmonic motion at any instant of time is





given by, y = 0.2sin2π(350t − 0.30). Calculate i) the amplitude, ii) wave velocity, iii) wave length,





iv) frequency.

1
Expert's answer
2022-02-23T08:32:45-0500

1)

"A=0.2~m,"

2)

"v=0.2\cdot 2\pi\cdot 350=440~\frac ms,"

3)

"\lambda=0.2\cdot 350=70~m,"

4)

"\nu={350}\cdot 2\pi=2.2~kHz."